OpenStudy (anonymous):

i need to solve this over the interval [0,2pi]. tan^2x=2secx-1

OpenStudy (anonymous):

tan^2 x - 2 sec x +1

OpenStudy (anonymous):

i simplified tan^2x into sec^2x-1. is that a good step?

OpenStudy (anonymous):

(sec^2 x -1) - 2 sec x +1 sec^2 x - 2 sec x

OpenStudy (anonymous):

sec (x)=u so substituting u^2-2u=0 We may complete the square by adding 1 to both side u^2 -2u+1=1 (u-1)(u-1)=1 u=2 sec (x)=2 so since 1/sec(x)= cos(x) cos(x)=1/2 x=pi/3

OpenStudy (anonymous):

wait what is u?

OpenStudy (anonymous):

I just define u for sec(x) for simplicity

OpenStudy (anonymous):

ok so is the work with sec(x)=u a continuation of the post you did before that?

OpenStudy (anonymous):

U is used in place of sec(x), just to make it easier to see

OpenStudy (anonymous):

no, I did not need to use U. I could have instead solve it using sec(x)

OpenStudy (anonymous):

Solved without using u sec(x)^2-2sec(x)=0 We may complete the square by adding 1 to both side sec(x)^2 -2sec(x)+1=1 (sec(x)-1)(sec(x)-1)=1 sec(x)=2 sec (x)=2 so since 1/sec(x)= cos(x) cos(x)=1/2 x=pi/3

OpenStudy (anonymous):

could i have used the quadratic formula to solve from where it was secx^2-2secx=0?

OpenStudy (anonymous):

and also how do you know you have to complete the square?

OpenStudy (anonymous):

Yes, but it is much easier

OpenStudy (anonymous):

u^2-2u=0 I could see that by adding 1 to left side, I would be able to factor easily. But if you add something one one side, same must be done on the other side

OpenStudy (anonymous):

thank you!